OpenStudy (batman19991):

All of the following have the same solution except _____. 50 = 5m b + 2 = 8 x - 6 = 4 w/5 = 2

OpenStudy (misssunshinexxoxo):

What looks different here? out of all of them? the symbol

OpenStudy (batman19991):

positive and negitives?

OpenStudy (misssunshinexxoxo):

All of these have 10 involved except for one

OpenStudy (misssunshinexxoxo):

b + 2 = 8 the solution is 6

OpenStudy (batman19991):

how though

OpenStudy (welshfella):

b + 2 = 8 subtract 2 from each side of the equation:- b + 2 - 2 = 8 - 2 b = 6

OpenStudy (batman19991):

positive and negitives?